### v1

@@ 2.2 Multi-reference distributionally robust optimization (MR-DRO) model
Consider a decision vector as $\bm{x}\in\mathbb{R}^{K}$ that resides in a feasible set $\mathbb{X}\subseteq\mathbb{R}^{K}$. A single-stage stochastic optimization model, as shown in (2), seeks an optimal solution of $\bm{x}$ that minimizes the expected value of a cost-related objective function $\ell(\bm{x},\bm{\xi})$, given by $\mathbb{E}^{\mathbb{P}}\left[\ell(\bm{x},\bm{\xi})\right]=\int_{\Xi}\ell(\bm{x},\bm{\xi})\mathbb{P}(d\bm{\xi})$. Here, $\Xi\subseteq\mathbb{R}^{M}$ represents the uncertainty set, and $\ell:\mathbb{R}^{K}\times\mathbb{R}^{M}\rightarrow\mathbb{R}$ is the objective function dependent on both the decision vector $\bm{x}$ and the uncertain parameters $\bm{\xi}$.
@@
We assume that $\mathbb{X}$ is a closed set and that the objective function $\ell(\bm{x},\bm{\xi})$ is lower semi-continuous in $\bm{x}$ for every fixed $\bm{\xi}\in\Xi$, and upper semi-continuous in $\bm{\xi}$ for every fixed $\bm{x}\in\mathbb{X}$. However, the unknown nature of the true distribution $\mathbb{P}$ prevents us from solving (2) directly. To address this issue, we formulate a multi-reference distributionally robust optimization (MR-DRO) problem:
@@
which includes all distributions $\mathbb{Q}$ on the space $\mathcal{M}(\Xi)$ such that $\mathbb{E}^{\mathbb{Q}}\left[\left\|\bm{\xi}\right\|\right]=\int_{\Xi}\left\|\xi\right\|\mathbb{Q}(d\bm{\xi})<\infty$ and their Wasserstein distance to ^$\hat{\mathbb{P}}_{HI}$ (obtained via (1)) is at most $\epsilon$ ($\epsilon\geq 0$). It can be viewed as a Wasserstein ball of radius $\epsilon$ centered on the empirical distribution ^$\hat{\mathbb{P}}_{HI}$. The Wasserstein distance is further defined as:
@@
[…]
The objective of this MR-DRO model (3) is to minimize the worst-case expectation of $\ell(\bm{x},\bm{\xi})$ in all possible distributions within the ambiguity set $\mathbb{B}_{\epsilon}(\hat{\mathbb{P}}_{HI})$. By carefully choosing the radius $\epsilon$, we can ensure that the true distribution $\mathbb{P}$ is contained within the ambiguity set with high probability. The inner maximization over probability distributions in Model (3) involves an optimization problem in an infinite-dimensional space, which may appear intractable. In Section 3 we will demonstrate that under reasonable assumptions for the uncertainty set $\Xi$ and the objective function $\ell(\bm{x},\bm{\xi})$, Model (3) can be reformulated into computationally tractable forms.
@@ 3 Tractable Reformulations of the MR-DRO Model
In this section, we demonstrate that the inner worst-case expectation problem in (3) over the Wasserstein ambiguity set (4), expressed as:
@@
can be reformulated as a finite convex program for many specific forms of the objective function $\ell(\bm{x},\bm{\xi})$. Without loss of generality, we assume that the objective function is defined as the point-wise maximum of $J$ elementary measurable functions $\ell_{j}:\mathbb{R}^{M}\times\mathbb{R}^{K}\rightarrow\mathbb{R}$, $j\in[J]$, so we have $\ell(\bm{x},\bm{\xi}):=\max_{j\in[J]}\left\{\ell_{j}(\bm{x},\bm{\xi})\right\}$. The following assumption on $\Xi$ and $\ell_{j}$ ensures tractability of the MR-DRO model. {assumption}[Convexity, Esfahani and Kuhn (2018), Assumption 4.1] The uncertainty set $\Xi\subseteq\mathbb{R}^{M}$ is convex and closed. Each $-\ell_{j}$ is proper (Charalambos D. Aliprantis 2006), convex, and lower semi-continuous in $\bm{\xi}$ for every fixed $\bm{x}\in\mathbb{X}$ and for all $j\in[J]$. In addition, $\ell_{j}$ is not identically $\infty$- ∞ in $\Xi$ for every fixed $\bm{x}\in\mathbb{X}$ and for all $j\in[J]$.
We then address the reduction of the inner worst-case expectation to a finite convex program in Section 3.1, and the reformulations for specific objective functions in Section 3.2.
@@ 3.1 Convex reduction
Given Assumption 6, then for any $\epsilon\geq 0$ the inner worst-case expectation (6) equals the optimal value of the finite convex program (7):
@@
Here, the conjugate of a function $f$ is defined as $f^{*}(\bm{v}):=\sup_{\bm{u}\in\text{dom }f}\langle\bm{v},\bm{u}\rangle-f(\bm{u})$. $\|\cdot\|_{*}$ is the dual norm of the norm used in the Wasserstein distance (5), which is defined through $\|\bm{v}\|_{*}:=\sup_{\|\bm{u}\|\leq 1}\langle\bm{v},\bm{u}\rangle$. The support function of $\Xi$ is defined as $\sigma_{\Xi}(\bm{u}):=\sup_{\bm{\xi}\in\Xi}\langle\bm{u},\bm{\xi}\rangle$.
A detailed proof of Theorem 3.1 is provided in Appendix A.1. This result forms the foundation for further reformulations under specific structural assumptions on the loss function.
3.2 Reformulations with regards to certain loss functions
@@
We present the proof of Theorem 3.2 in Appendix A.2. This reformulation provides a computationally efficient framework for handling piecewise affine loss functions.
Note that if we use $L_{1}$-norm in the definition of Wasserstein distance (5), the dual norm here becomes $L_{\infty}$-norm. Furthermore, if the feasible set $\mathbb{X}\subseteq\mathbb{R}^{K}$ is a set of linear constraints with regard to $\bm{x}$, we end up with a linear programming (LP) reformulation (8).
@@
[Separable affine objective functions] Suppose that the uncertain parameters can be separated as $\bm{\xi}=(\bm{\xi}_{1},\ldots,\bm{\xi}_{N})$, and assume that $\bm{\xi}_{n}\in\Xi_{n}$, where $\Xi_{n}\subseteq\mathbb{R}^{M}$ is nonempty and closed for any $n\in[N]$. The objective function is additively separable with respect to the structure of $\bm{\xi}$, so we have $\ell(\bm{x},\bm{\xi}):=\sum_{n=1}^{N}\max_{j\in[J]}\left\{\ell_{nj}(\bm{x},\bm{\xi}_{n})\right\}$.
[…]
@@
Under Assumptions 3.2 and 3.2, we have $\Xi_{n}=\left\{\bm{\xi}_{n}\in\mathbb{R}^{M}:\bm{C}_{n}\bm{\xi}_{n}\leq\bm{g}_{n}\right\}$ where $\bm{C}_{n}$ is a matrix and $\bm{g}_{n}$ is a vector of appropriate dimension for all $n\in[N]$. In addition, for every fixed $\bm{x}\in\mathbb{X}$, each elementary function $\ell_{nj}(\bm{x},\bm{\xi}_{n})=\langle\bm{a}_{nj},\bm{\xi}\rangle+b_{nj}$ is an affine function with regard to $\bm{\xi}_{n}$, where $\bm{a}_{nj}\in\mathbb{R}^{M}$ and $b_{nj}\in\mathbb{R}$ are linear in $\bm{x}$ for all $n\in[N]$ and $j\in[J]$. Then for any $\epsilon\geq 0$, the MR-DRO model (3) can be reformulated as:
$\inf_{\bm{x},\lambda,s_{hik},\bm{\gamma}_{hijk}}\quad$ $\lambda\epsilon+\sum^{H}_{h=1}\sum^{I}_{i=1}\sum^{N}_{n=1}\frac{t^{(I)}_{h}}{I}\cdot s_{hik}$ (9a)
@@
$b_{nj}+\langle\bm{a}_{nj},\hat{\bm{\xi}}^{(i)}_{h,n}\rangle+\langle\bm{\gamma}_{hijn},\bm{g}_{n}-\bm{C}_{n}\hat{\bm{\xi}}^{(i)}_{h,n}\rangle\leq s_{hik},$ $\forall h\in[H],i\in[I],j\in[J],n\in[N],$ (9c)
@@
The proof of Theorem 3.3, detailed in Appendix A.3, builds on the framework established in Theorem 3.2 by exploiting the separability of the loss function. This allows for computational simplifications in scenarios where the uncertain parameters are structured as separable components.
@@ 4 Dynamic Trust Update Process
We now design a dynamic trust update mechanism to simulate decision makers’ trust in various sources evolving over events as more data become available. This approach serves two primary purposes. First, it aims to improve decision accuracy and reliability by leveraging the discrete distribution ^$\hat{\mathbb{P}}_{HI}$, derived by combining nonparametric data using trust as probabilistic weights, which forms the center of the Wasserstein ambiguity set in our MR-DRO model. By updating trust values, the model can mitigate the impact of poor data quality by reducing the weights assigned to less reliable sources. Second, updating trust values based on the performance of each source ensures that the model reflects real-world trust dynamics, where a decision maker’s trusts in multiple sources are adjusted according to their track records. Increasing trust values in sources that provide accurate predictions while decreasing trust in those that do not mirror natural human behavior, thereby enhancing the model’s adaptability.
An approach to simulate trust dynamics is to utilize actual losses from each event as an evaluation metric (Guo et al. 2024). Through the trial-and-error method, one can estimate the partial derivative of the actual losses over the trust in one source and then update it accordingly. A negative partial derivative of actual losses with respect to the current trust in one data source indicates that increasing trust in this source will result in decreased losses, and therefore in the next iteration, we will increase the trust. A positive partial derivative works in the opposite way. This is a handy method for the parametric data fusion way to construct the ambiguity set, and can also be implemented even when we do not have historical prediction errors from each source.
In this paper, we propose an error-based trust update method by examining the historical data misalignment between the predictions from each source and the true realization values of the uncertainty for past events in Section 4.1. We also generalize two algorithms: the static expert algorithm and the variable share algorithm in Herbster and Warmuth (1998) to update trust in Sections 4.2 and 4.3, respectively. Furthermore, in Section 4.4, we introduce the concept of probability dominance and show how to identify sources with dominant trust under certain assumptions.
@@ 4.1 Min-max error trust update
We decompose each event as a 4-step procedure, shown in Figure 1.
@@
At the beginning of event $i$, we have trust $t^{(i-1)}_{h}$ in source $h$, for all $h\in[H]$. Then we update the ambiguity set $\mathbb{B}_{\epsilon}(\hat{\mathbb{P}}_{HI})$ and solve MR-DRO to attain an optimal solution $\bm{x}^{*}(\bm{t}^{(i-1)})$. After event $i$ ends, we obtain realization $\bm{\xi}^{(i)}_{\textrm{true}}$ of the uncertain parameter $\bm{\xi}$. As events occur independently, each $\bm{\xi}^{(i)}_{\textrm{true}}$ is independent of the others for all $i\in[I]$.
@@
[…]
@@
Revisiting Example 2.1 discussed in Section 2.1, we now illustrate how we apply the min-max error trust update algorithm within the same context. Using the empirical distribution ^$\hat{\mathbb{P}}_{HI}$, where $H=2$ and $I=2$, we construct the ambiguity set and solve the MR-DRO model to obtain an optimal solution $x^{*}(\bm{t}^{(2)})$. We would like to update the trust to $\bm{t}^{(3)}$ for the upcoming event. Suppose that, after the current event, we obtain the realization as $\xi^{(3)}_{\text{true}}=8$. Therefore, we obtain the prediction errors for the current event from each source as $\Delta\xi^{(3)}_{1}=-2$ and $\Delta\xi^{(3)}_{2}=1$. Suppose that we use $L_{1}$-norm as our evaluation criterion. Since $|\Delta\xi^{(3)}_{1}|>|\Delta\xi^{(3)}_{2}|$, this indicates that the prediction of source $1$’s was less accurate than the prediction of source $2$’s for the current event. As a result, we decrease $t^{(3)}_{1}$ by a defined step size $\Delta t$ and correspondingly increase $t^{(3)}_{2}$ by $\Delta t$ for the next event.

### v2

@@ 2.2 Multi-reference distributionally robust optimization (MR-DRO) model
Consider decision $\bm{x}\in\mathbb{R}^{K}$ that resides in a feasible set $\mathbb{X}\subseteq\mathbb{R}^{K}$. A stochastic optimization model:
@@
seeks an optimal solution of $\bm{x}$ that minimizes the expected value of a cost-related objective function $\ell(\bm{x},\bm{\xi})$, given by $\mathbb{E}^{\mathbb{P}}\left[\ell(\bm{x},\bm{\xi})\right]=\int_{\Xi}\ell(\bm{x},\bm{\xi})\mathbb{P}(d\bm{\xi})$. Here, $\Xi\subseteq\mathbb{R}^{M}$ represents the uncertainty set, and $\ell:\mathbb{R}^{K}\times\mathbb{R}^{M}\rightarrow\mathbb{R}$ is the objective function dependent on both the decision vector $\bm{x}$ and the uncertain parameters $\bm{\xi}$.
By assumption, $\mathbb{X}$ is a closed set and the objective function $\ell(\bm{x},\bm{\xi})$ is lower semi-continuous in $\bm{x}$ for every fixed $\bm{\xi}\in\Xi$, and upper semi-continuous in $\bm{\xi}$ for every fixed $\bm{x}\in\mathbb{X}$. However, the unknown nature of the true distribution $\mathbb{P}$ prevents us from solving (2) directly. To address this issue, we formulate a multi-reference distributionally robust optimization (MR-DRO) problem:
@@
which includes all distributions $\mathbb{Q}$ on the space $\mathcal{M}(\Xi)$ such that $\mathbb{E}^{\mathbb{Q}}\left[\left\|\bm{\xi}\right\|\right]=\int_{\Xi}\left\|\xi\right\|\mathbb{Q}(d\bm{\xi})<\infty$ and their Wasserstein distance to $\hat{\mathbb{P}}_{HI}$ (obtained via (1)) is at most $\epsilon$ ($\epsilon\geq 0$). It can be viewed as a Wasserstein ball of radius $\epsilon$ centered on the empirical distribution $\hat{\mathbb{P}}_{HI}$. The Wasserstein distance is further defined as:
@@
where $\|\cdot\|$ represents an arbitrary norm on $\mathbb{R}^{M}$, and $\Pi$ is a joint distribution of $\bm{\xi}$ and $\bm{\xi}^{{}^{\prime}}$ with marginal distributions $\mathbb{Q}$ and $\hat{\mathbb{P}}_{HI}$, respectively.
The objective of this MR-DRO model (3) is to minimize the worst-case expectation of $\ell(\bm{x},\bm{\xi})$ in all possible distributions within the ambiguity set $\mathbb{B}_{\epsilon}(\hat{\mathbb{P}}_{HI})$. The inner maximization over probability distributions in Model (3) involves an optimization problem in an infinite-dimensional space, which may appear intractable. In Section 3, we demonstrate that under reasonable assumptions for the uncertainty set $\Xi$ and the objective function $\ell(\bm{x},\bm{\xi})$, Model (3) can be reformulated into computationally tractable forms.
@@ 3 Tractable Reformulations of the MR-DRO Model
We show that the inner worst-case expectation problem in (3) over the Wasserstein ambiguity set (4), expressed as:
@@
can be reformulated as a finite convex program for many specific forms of the objective function $\ell(\bm{x},\bm{\xi})$. Without loss of generality, we assume that the objective function is defined as the point-wise maximum of $J$ elementary measurable functions $\ell_{j}:\mathbb{R}^{M}\times\mathbb{R}^{K}\rightarrow\mathbb{R}$, $j\in[J]$, and therefore, $\ell(\bm{x},\bm{\xi}):=\max_{j\in[J]}\left\{\ell_{j}(\bm{x},\bm{\xi})\right\}$. The following assumption on $\Xi$ and $\ell_{j}$ ensures tractability of the MR-DRO model. {assumption}[Convexity, \@BBOPcitet\@BAP\@BBNEsfahani and Kuhn (2018)\@BBCP, Assumption 4.1] The uncertainty set $\Xi\subseteq\mathbb{R}^{M}$ is convex and closed. Each $\ell_{j}$ is proper \@BBOPcitep\@BAP\@BBN(Charalambos D. Aliprantis 2006)\@BBCP, convex, and lower semi-continuous in $\bm{\xi}$ for every fixed $\bm{x}\in\mathbb{X}$ and for all $j\in[J]$. In addition, $\ell_{j}$ is not identically -$\infty$ in $\Xi$ for every fixed $\bm{x}\in\mathbb{X}$ and for all $j\in[J]$.
We address the reduction of the inner worst-case expectation to a finite convex program in Section 3.1, and the reformulations for specific objective functions in Section 3.2.
@@ 3.1 Convex reduction
Given Assumption 6, then for any $\epsilon\geq 0$ the inner worst-case expectation (6) equals the optimal objective value of the following convex program (7):
@@
Here, the conjugate of a function $f$ is defined as $f^{*}(\bm{v}):=\sup_{\bm{u}\in\text{dom }f}\langle\bm{v},\bm{u}\rangle-f(\bm{u})$; notation $\|\cdot\|_{*}$ is the dual norm of the norm used in the Wasserstein distance (5), which is defined through $\|\bm{v}\|_{*}:=\sup_{\|\bm{u}\|\leq 1}\langle\bm{v},\bm{u}\rangle$. The support function of $\Xi$ is defined as $\sigma_{\Xi}(\bm{u}):=\sup_{\bm{\xi}\in\Xi}\langle\bm{u},\bm{\xi}\rangle$.
A detailed proof of Theorem 3.1 is provided in Supplement Material Section A.1. This result forms the foundation for further reformulations under specific structural assumptions on the loss function.
3.2 Reformulations for certain loss functions
@@
We present the detailed proof of Theorem 3.2 in Supplement Material Section A.2. This reformulation provides a computationally efficient framework for handling piecewise affine loss functions.
Note that if we use $L_{1}$-norm in the definition of Wasserstein distance (5), the dual norm here becomes $L_{\infty}$-norm. Furthermore, if the feasible set $\mathbb{X}\subseteq\mathbb{R}^{K}$ is a set of linear constraints with regard to $\bm{x}$, we have a linear programming (LP) reformulation (8).
@@
[Separable affine objective functions] Suppose that the uncertain parameters can be separated as $\bm{\xi}=(\bm{\xi}_{1},\ldots,\bm{\xi}_{N})$, and assume that $\bm{\xi}_{n}\in\Xi_{n}$, where $\Xi_{n}\subseteq\mathbb{R}^{M}$ is nonempty and closed for any $n\in[N]$. The objective function is additively separable with respect to the structure of $\bm{\xi}$, i.e., $\ell(\bm{x},\bm{\xi}):=\sum_{n=1}^{N}\max_{j\in[J]}\left\{\ell_{nj}(\bm{x},\bm{\xi}_{n})\right\}$.
Note that such loss functions commonly appear, for example, in multi-item newsvendor problems and production planning problems under uncertain demands. We define $\|\bm{\xi}\|_{N}=\sum_{n=1}^{N}\|\bm{\xi}_{n}\|$ to measure the overall uncertainty across $N$ and also require the $\|\cdot\|$ on $\mathbb{R}^{M}$ to be the same norm we used in the definition of the Wasserstein metric (5).
@@
Given Assumptions 3.2 and 3.2, we have $\Xi_{n}=\left\{\bm{\xi}_{n}\in\mathbb{R}^{M}:\bm{C}_{n}\bm{\xi}_{n}\leq\bm{g}_{n}\right\}$ where $\bm{C}_{n}$ is a matrix and $\bm{g}_{n}$ is a vector of appropriate dimension for all $n\in[N]$. In addition, for every fixed $\bm{x}\in\mathbb{X}$, each elementary function $\ell_{nj}(\bm{x},\bm{\xi}_{n})=\langle\bm{a}_{nj},\bm{\xi}\rangle+b_{nj}$ is an affine function with regard to $\bm{\xi}_{n}$, where $\bm{a}_{nj}\in\mathbb{R}^{M}$ and $b_{nj}\in\mathbb{R}$ are linear in $\bm{x}$ for all $n\in[N]$ and $j\in[J]$. Then for any $\epsilon\geq 0$, the MR-DRO model (3) can be reformulated as:
$\inf_{\bm{x},\lambda,s_{hin},\bm{\gamma}_{hijn}}\quad$ $\lambda\epsilon+\sum^{H}_{h=1}\sum^{I}_{i=1}\sum^{N}_{n=1}\frac{t^{(I)}_{h}}{I}\cdot s_{hin}$ (9a)
@@
$b_{nj}+\langle\bm{a}_{nj},\hat{\bm{\xi}}^{(i)}_{h,n}\rangle+\langle\bm{\gamma}_{hijn},\bm{g}_{n}-\bm{C}_{n}\hat{\bm{\xi}}^{(i)}_{h,n}\rangle\leq s_{hin},$ $\forall h\in[H],i\in[I],j\in[J],n\in[N],$ (9c)
@@
The proof of Theorem 3.3, detailed in Supplement Material Section A.3, builds on the framework established in Theorem 3.2 by exploiting the separability of the loss function. This allows for computational simplifications in cases where the uncertain parameters are structured as separable components.
@@ 4 Dynamic Trust Update Process
We now introduce a dynamic trust update mechanism to capture how a decision maker’s trust in multiple data sources evolves over time as new information becomes available. This mechanism serves two primary objectives. First, it enhances decision accuracy and robustness by leveraging the discrete distribution $\hat{\mathbb{P}}_{HI}$—constructed through nonparametric data fusion with trust values acting as probabilistic weights—which serves as the center of the Wasserstein ambiguity set in the MR-DRO model. By dynamically updating trust levels, the framework reduces the influence of low-quality or unreliable data sources, thereby mitigating adverse effects arising from inaccurate information. Second, the adaptive adjustment of trust values based on observed performance aligns the model with realistic decision-making behavior, where confidence in different sources evolves according to their historical reliability. Increasing trust in sources that consistently yield accurate predictions while decreasing trust in poorly performing ones mirrors natural human learning processes and improves the model’s adaptability to changing environments.
One approach to modeling trust dynamics is to use realized losses from each event as a performance metric \@BBOPcitep\@BAP\@BBN(Guo et al. 2024)\@BBCP. Under a trial-and-error framework, one may approximate the partial derivative of realized losses with respect to the trust assigned to a particular source and update trust accordingly. A negative partial derivative indicates that increasing trust in that source reduces losses, suggesting that its weight should be increased in subsequent iterations, whereas a positive derivative implies the opposite adjustment. This gradient-based approach is particularly convenient in parametric data fusion settings for constructing ambiguity sets and remains applicable even when historical prediction error data are unavailable.
In this paper, we instead propose an error-based trust update scheme that evaluates historical misalignment between each source’s predictions and the realized uncertainty outcomes, as detailed in Section 4.1. Building on this idea, we generalize two classical online learning methods—the static expert algorithm and the variable-share algorithm of \@BBOPcitet\@BAP\@BBNHerbster and Warmuth (1998)\@BBCP—to update trust values, as presented in Sections 4.2 and 4.3. Finally, Section 4.4 introduces the concept of probability dominance and establishes conditions under which sources with dominant trust can be identified.
@@ 4.1 Min-max error trust update
We decompose each event as a four-step procedure, shown in Figure 1.
@@
At the beginning of event $i$, we have trust $t^{(i-1)}_{h}$ for each data source $h$, for all $h\in[H]$. Then we update the ambiguity set $\mathbb{B}_{\epsilon}(\hat{\mathbb{P}}_{HI})$ and solve MR-DRO to attain an optimal solution $\bm{x}^{*}(\bm{t}^{(i-1)})$. After event $i$ ends, we obtain realization $\bm{\xi}^{(i)}_{\textrm{true}}$ of the uncertain parameter $\bm{\xi}$. As events occur independently, each $\bm{\xi}^{(i)}_{\textrm{true}}$ is independent of the others for all $i\in[I]$.
@@
In this method, the trust update is driven by the prediction error from each source $\Delta\bm{\xi}^{(i)}_{h}$ in the most recent event, evaluated from a norm $\|\cdot\|$ on $\mathbb{R}^{M}$. For each update, we only focus on the extreme cases: a source with the minimum norm of prediction error will experience an increase in its trust value, while we will decrease the trust in the source with the maximum norm. To ensure stability, we initialize $t^{(0)}_{h}=Q/R$ for all $h\in[H]$ where $Q,R$ are positive integers with the highest common divisor $1$. The step size for the update is defined as $\Delta t=1/qR$, where $q\in\mathbb{N}_{+}$.
@@
[…]